OpenStudy (anonymous):

1.What is 7* 4-2000+30*3= you will get a medal any one

OpenStudy (saifoo.khan):

PEDMAS! Heard of it?

OpenStudy (anonymous):

Parenthesis Exponets Multiplacation Division Addition Subtraction

OpenStudy (kendricklamar2014):

Using PEMDAS: 7 * 4 = 28 30 * 3 = 90

OpenStudy (kendricklamar2014):

2,000 + 90 = 2,090

OpenStudy (kendricklamar2014):

2,090 - 28 = 2,062

OpenStudy (anonymous):

are you sure

OpenStudy (anonymous):

your medal

OpenStudy (kendricklamar2014):

Yes I am sure. Next time, use PEMDAS :)